OpenStudy (anonymous):

find f(g(x)) and g(f(x)) and determine whether the pair of functions f and g are inverse f(x)=3x-6 and g(x) x+3/6 a) f(g(x))= b)g(f(x))=

OpenStudy (anonymous):

is it \[f(x)=3x-6\] and \[g(x)=\frac{x+6}{3}\]? if so , they are definitely inverses

OpenStudy (anonymous):

\(f\) says multiply by 3, subtract 6 \(g\) says add 6, divide by 3 inverse operations in reverse order

OpenStudy (alexwee123):

to find inverse you just switch the x and y values around then isolate y :o

OpenStudy (mertsj):

The directions say to find the f(g(x)) and the g(f(x)). If you get x for both of those, that shows the functions are inverse functions.

OpenStudy (anonymous):

\[f(g(x))=f(\frac{x+6}{3})=3(\frac{x+6}{3})-6=\text{some algebra}=x\]

OpenStudy (anonymous):

If f(g(x)) = g(f(x)), then they are inverses.

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

Yeah. My teacher told me.

OpenStudy (mertsj):

Only if they are both equal to x

OpenStudy (anonymous):

\[f(x)=2x, g(x)=3x\] \[f(g(x))=g(f(x))=6x\] i sincerely hope your teacher did not tell you that

OpenStudy (anonymous):

@avatar2012 you job is to compute \[f(g(x))\] and \[g(f(x))\] i wrote the first one. you will get \(x\) back for the second one as well, so they are inverses for sure

OpenStudy (anonymous):

If they are truly inverse, such as f(x) = 2x + 6 and g(x) = (x - 6)/2, then f(g(x)) = g(f(x)) f(g(x)) = (2)(x - 6)/2 + 6 = x - 6 + 6 = x g(f(x)) = (2x + 6 - 6)/2 = 2x/2 = x See? x = x

OpenStudy (anonymous):

These are not inverse because: y = 3x - 6 x = 3y - 6 x + 6 = 3y (x + 6)/3 = y (x + 3)/6 ≠ (x + 6)/3

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

thanks

OpenStudy (mertsj):

They are inverse functions!!

OpenStudy (anonymous):

f(x)=3x-6 and g(x) x+3/6 are inverse functions? Because satellite73 wrote it wrong.

OpenStudy (anonymous):

y = 3x - 6 x = 3y - 6 x + 6 = 3y (x + 6)/3 = y (x + 3)/6 ≠ (x + 6)/3? That's what I did.